OpenStudy (anonymous):

The statement below is justified by the _________. 4 + 4x = 4x + 4 commutative property of addition associative property of addition addition property of equality additive identity property

jimthompson5910 (jim_thompson5910):

check out page 2 of this pdf http://www.cengage.com/resource_uploads/downloads/1133963021_375161.pdf

OpenStudy (anonymous):

I don't get it

jimthompson5910 (jim_thompson5910):

did you read the properties on page 2?

OpenStudy (anonymous):

Yeah I just don't get how to do it AT ALL

jimthompson5910 (jim_thompson5910):

look for the rule that says a+b = b+a

jimthompson5910 (jim_thompson5910):

this rule effectively tells us that we can add two numbers in any order example 2+3 = 5 3+2 = 5 so...2+3 = 3+2